The function a*b is defined as a*b = a^2b-a+b.What is the value of 2*(1*(2*(1*(2*1))))?
Answer by stanbon(75887)

The function a*b is defined as a*b = a^2b-a+b.
What is the value of 2*(1*(2*(1*(2*1))))?
-------
2*(1*(2*(1*(2*1)))) = 2*(1*(2*(2^2*1-2+1))))
----------
= 2*(1*(2*3)))
----
= 2*(1*(2^2*3-2+3)))
-----
= 2*(1*11)
------
= 2*(1^2*11-1+11)
------
= 2*21
-----
= 2^2*21-2+21
-----
= 84-2+21
= 103
------------
Cheers,
Stan H.
-----------